### v2

@@ 7 Conclusion
This paper presented BernGraph, a novel approach that handles sparse sequences in EHRs by transforming the problematic 0,1 binary event outcomes to continuous Bernoulli means. We took a statistical perspective by viewing the EHR data as a sample from the greater population. In this manner, we modeled the nodes and edges of GNN using the Bernoulli distribution. Extensive results showed that our method attained the SOTA and outperformed all baselines by a large margin, and some of the baselines learned from significantly more information resources, such as the visiting data and drug-drug interactions, while ours learned solely from the binary data. Ablation studies verified the importance of the Bernoulli transformation of EHR binary data. By replacing only edges (conditional probability) and keeping nodes (Bernoulli mean) fixed, we observed largely reduced improvement compared to the proposed method. On the other hand, assigning Bernoulli means to the nodes resulted in significantly better performance, even with random edge values. These findings confirm that BernGraph is a simple and effective model, highlighting its potential for integration into more sophisticated systems.
Appendix 0.A Source Code
Appendix 0.B Preprocessing
0.B.1 Data Profile
@@
0.B.2 Medication Data Preprocessing
@@
0.B.3 Diagnosis Data Preprocessing
@@
0.B.4 Procedure Data Preprocessing
@@ 0.B.5 Data Integration
Appendix 0.C Model Parameters
@@ Appendix 0.D Metrics
Log Likelihood Ratio (LLR): The likelihood ratio is defined as the ratio of two key values: The maximum value of the likelihood function is calculated within the subspace defined by the hypothesis. And the maximum value of the likelihood function, calculated across the entire parameter space [8].
@@
In our case, when calculating the similarity measure between a medical event and the recommendation of a specific drug, we calculated the LLR value by the counts that co-occurred between them. The average LLR for each medical event with all recommended drugs was calculated as the properties of the medical event node [8].
Appendix 0.E Implementation Details
@@
Appendix 0.F More Results
Figure 5 shows another visualization of features learned by GNN from patients’ EHR data. This time we peak another set of medical events on other two different patients.
Figure 5: Another visualization of features learned by GNN from patients’ EHR data. Each node represents a different event, and edges represent relationships between events. Thicker edges indicate a stronger association between two events. Nodes with higher feature values have been highlighted.

### v3

@@ 7 Conclusion
This paper introduces BernGraph, a novel approach that transforms sparse 0/1 binary event sequences in EHRs into continuous Bernoulli means. Viewing EHR data as samples from a population, we modeled GNN nodes and edges using the Bernoulli distribution. Our method outperformed all baselines, including those using more extensive data like visit records and drug-drug interactions. Ablation studies confirmed the importance of Bernoulli transformation, with node-based Bernoulli means contributing significantly to performance, even with random edges. This work presents a promising direction for handling sparse binary sequences in EHRs, potentially improving various healthcare prediction tasks.
Appendix A Preprocessing
A.1 Data Profile
@@
A.2 Medication Data Preprocessing
@@
A.3 Diagnosis Data Preprocessing
@@
A.4 Procedure Data Preprocessing
@@ A.5 Data Integration
Appendix B Model Parameters
@@ Appendix C Metrics
Log Likelihood Ratio (LLR): The likelihood ratio is defined as the ratio of two key values: The maximum value of the likelihood function is calculated within the subspace defined by the hypothesis. And the maximum value of the likelihood function, calculated across the entire parameter space (Dunning, 1994).
@@
In our case, when calculating the similarity measure between a medical event and the recommendation of a specific drug, we calculated the LLR value by the counts that co-occurred between them. The average LLR for each medical event with all recommended drugs was calculated as the properties of the medical event node (Dunning, 1994).
Appendix D Implementation Details